OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative. Help Please!!!!

OpenStudy (anonymous):

just differintiate the function.. \[V(t)=-9 \rightarrow V(1) =-9\]

OpenStudy (anonymous):

i'm asking this because i'm really clueless when it comes to differentiating. Can you please explain how to?

OpenStudy (anonymous):

when differentiating any function use the formula : \[(ax ^{n} )\prime = a*n x ^{n-1} \] such that: a is constant (-9 in the case provided) and n is the power of x also the derivative of the constants is always 0 I think that's all u need to know at the moment

OpenStudy (anonymous):

I am sorry not any function only polynomials and few others like \[\sqrt{x} , \frac{ 1 }{ x }\]